OpenStudy (anonymous):

Find the length of the arc on a circle of radius 7 ft intercepted by a central angle 248 ◦. Answer in units of ft

OpenStudy (anonymous):

\[\pi d \frac{248{}^{\circ}}{360{}^{\circ}}=\pi 14 \frac{248{}^{\circ}}{360{}^{\circ}}=\frac{434 \pi }{45}=30.2989 \]

OpenStudy (anonymous):

Thank you for the medal(s)